In construction, floor space must be given for staircases. If the second floor is 3.6 meters above the first floor and a contractor is using the standard step pattern of 28 cm of tread for 19 cm of rise then how many steps are needed to get from the first to the second floor and how much linear distance will need to be used for the staircase?
Found 2 solutions by stanbon, Paul:
Answer by stanbon(48558)

In construction, floor space must be given for staircases. If the second floor is 3.6 meters above the first floor and a contractor is using the standard step pattern of 28 cm of tread for 19 cm of rise then how many steps are needed to get from the first to the second floor and how much linear distance will need to be used for the staircase?
3.6 meters = 360 cm of rise
360/19=18.95
You need 19 steps to get a rise of 3.6 meters
19 steps cover a linear distance of 19*28 cm = 5.32 meters
Cheers,
Stan H.

Answer by Paul(988)

3.6 m = 360 cm
360/19=19 risers
19 steps are needed
19*28=532
Hence, 19 steps cover 532 cm.
Paul.